If (x,12,13) is a Pythagoreantriple, then x = : If (x,12,13) is a Pythagoreantriple, then x =
Answer by stanbon(18724)

If (x,12,13) is a Pythagorean triple, then x =
------------
13^2 = 12^2 + x^2
x^2 = 13^2 - 12^2
x^2 = 169 - 144
x^2 = 25
x = 5
==========
Cheers,
Stan H.
Question 142488: If (x,12,13) is a Pythagoreantriple, then x = : If (x,12,13) is a Pythagoreantriple, then x =
Answer by benni1013(141)

The answer is 5. A pythagorean triple a^2+b^2=c^2 checks out very nicely.